The manager of a small business hires a consultant to help determine how the manager can improve the employees' moral. Currently there are 8 employees who like their job (have high moral) and 6 who do not like their job (have low moral). The consultant randomly selects 3 workers to interview. What is the probability the consultant selects three employees who like their job?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)

Why did you post the same question twice?

Answer by stanbon(75887)

The manager of a small business hires a consultant to help determine how the manager can improve the employees' morale. Currently there are 8 employees who like their job (have high morale) and 6 who do not like their job (have low morale). The consultant randomly selects 3 workers to interview. What is the probability the consultant selects three employees who like their job?
-------------------------
# of ways to succeed: 8C3 = (8*7*6)/(1*2*3) = 56
---
# of random groups of three: 14C3 = 14*13*12/1*2*3 = 364
------
P(3 who like their job) = 56/364 = 0.1538
===========================
Cheers,
Stan H.